OpenStudy (nerdystudieshelp):

Copper has a density of 8.96 g/cm^3. If 75.0 g of copper is added to 50.0 mL of water in a graduated cylinder, to what volume reading will the water level in the cylinder rise?

OpenStudy (aveline):

\[\frac{ 75.0g}{ 1 }\times \frac{ cm ^{3} }{ 8.96g }=?\]

OpenStudy (aveline):

@NerdyStudiesHelp

OpenStudy (zale101):

Please post your attempt to this problem.